OpenStudy (anonymous):

find the circumference and area of a circle with the diameter 6 inches.use 3.14 for pi

OpenStudy (anonymous):

whats the formula for area?

OpenStudy (anonymous):

Circle Area = π × r2 Circumference = 2 × π × r r = radius

OpenStudy (anonymous):

is that correct